OpenStudy (anonymous):

1. How is the decimal number 6 written in binary? 100 110 101 111

OpenStudy (jamesj):

10 in binary is 2 + 0 = 2 100 in binary is 4 + 0 + 0 = 4 So what do you think 6 is in binary?

OpenStudy (anonymous):

Uhm... 111?

OpenStudy (anonymous):

idk D:

OpenStudy (jamesj):

By definition of binary, 111 = 1 * 4 + 1 * 2 + 1 * 1 = 7 So 111 is not right

OpenStudy (anonymous):

Any conversion from decimal to binary is to keep dividing the the decimal number by 2 and push the remainder into a stack, when you reach 0 as the quotient it's time to pop the stack and furnish the result.

OpenStudy (anonymous):

110?

OpenStudy (jamesj):

Writing a subscript of 2 to denote the number is binary, \[ 110_2 = 1*4 + 1*2 + 0*1 = 6\] so yes

OpenStudy (anonymous):